OpenStudy (anonymous):

Giving Medals ((&)) Fan The table shows the solution to the equation |2x - 3| - 1 = 2: Step 1 |2x - 3| = 2 + 1 Step 2 |2x - 3| = 3 Step 3 2x - 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? A) Step 1 B) Step 3 C)Step 5 D)Solution is correct

OpenStudy (help!!!!):

B. Step three is incorrect because it is suppose to be 2x-3=3 OR 2x-3=-3

OpenStudy (anonymous):

Okay can you help me w/ 1 more ?

OpenStudy (anonymous):

Solve -4(x + 10) - 6 = -3(x - 2). A)-40 B)-46 C)-52 D)52

OpenStudy (anonymous):

@HELP!!!!